The equation for the instantaneous voltage across a discharging capacitor is given by, where is the initial voltage and is the time constant of the circuit

The tasks are to:

A) draw a graph of voltage against time for and, between and.
B) calculate the gradient at and
C) differentiate and calculate the value of at and
D) compare your answers with part b and part c
E) calculate the second derivative of the instantaneous voltage
1
Expert's answer
2020-12-23T04:32:44-0500

a)


b) "v=V_0\times e^{\frac{-t}{\tau}}" ;

1) "\frac{\partial v}{\partial V_0}= e^{\frac{-t}{\tau}}" ;

2)"\frac{\partial v}{\partial t}= -\frac{V_0}{\tau}\times e^{\frac{-t}{\tau}}" ;

3) "\frac{\partial v}{\partial t}= -V_0\times t \times e^{\frac{-t}{\tau}}" ;

c) "v^{\prime}=-6\times e^{-\frac{t}{2}}"

"v^{\prime}(2)=-2.21" ;

"v^{\prime}(4)=-0.81" ;

d) Derivative equal to the gradient from t

e) "v=\frac{V_0}{\tau^2}\times e^{-\frac{t}{\tau}}"